OpenStudy (anonymous):

y = −x^2 + 4 y = 2x + 1 How would you solve by elimination or substitution?

rvc (rvc):

we have : \[\rm y=-x^2+4~ and ~y=2x+1\]

OpenStudy (anonymous):

yes, I don't get how you would solve them though

rvc (rvc):

so we have two values for y

rvc (rvc):

we can put any of those

OpenStudy (anonymous):

like \[-x^2+4=2x+1\]

rvc (rvc):

yes

rvc (rvc):

now solve further :)

OpenStudy (anonymous):

but how would you subract them. Thats what im stuck on.

rvc (rvc):

here it is : \[\rm 2x+1+x^2-4\]

OpenStudy (anonymous):

that is what y equals then right?

rvc (rvc):

no since when we solve the equation we are left only with x. then we will put the value of x in any one of the original equation to find the value of y

OpenStudy (anonymous):

Im really stuck then .. how would you find what x and y are?

rvc (rvc):

here it is : \[\sf x^2+2x-3=0\]

rvc (rvc):

we have to solve the quadratic equation

OpenStudy (anonymous):

I get x = 1, -3

rvc (rvc):

yep now put one value at a time in either : y = 2x+1 or y=-x^2+4

OpenStudy (anonymous):

Thank you :)

rvc (rvc):

Thank you All the best :)